THE CO-DIMER: NEW STATES AND TUNNELING SPLITTING

\underline{L. SURIN}, D. FOURZIKOV, and G. WINNEWISSER, I. Physikalisches Institut, University of Cologne, Cologne, D 50937, Germany; B. DUMESH, V. PANFILOV, Insitute of Spectroscopy, 142190 Troitsk, Moscow region, Russia.

New extensive millimeter-wave measurements of the ^{1}^{2}C^{1}^{6}O dimer have been made and more than 250 new spectral transitions have been observed in the frequency range of 80 - 135 GHz. Joint analysis of these and previous millimeter-wave data yielded in determination and precise location of 34 new energy levels of A^{+}-symmetry and 21 levels of A^{−}-symmetry. Some of them belong to already known stacks and others make up 9 new stacks of the dimer. These new energy levels are located at energies from 8 to 18 cm^{−}^{1} and in a free rotation limit they correspond to the states with (j_{1}, j_{2}, K) = (1, 1, 0), (1, 1, 1), (1, 1, 2) and (2, 0, 2), where j is a rotational quantum numbers of the CO monomer and K is a projection of the total angular momentum on the intermolecular axis. One newly observed state of A^{+}-symmetry originating at 12 cm^{−}^{1} has K = 0 and might be tentatively attributed to the lower tunneling component of the stretching vibration of the CO-dimer. The tunneling splitting for many different states of two isotopic modifications of the dimer, (^{1}^{2}C^{1}^{6}O)_{2} and (^{1}^{3}C^{1}^{6}O)_{2} was determined, and its dependence on J-, K- values and on isotopic mass was studied. For some states the tunneling splitting increases in (^{1}^{3}C^{1}^{6}O)_{2} as compared to (^{1}^{2}C^{1}^{6}O)_{2}. Possible explanations of this anomalous behavior will be discussed.